OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I'll fan and give a medal!!

OpenStudy (abbycross167):

The width, w, of a rectangular rug is 4 less than its length, L . If the perimeter of the rug is 120 square feet, which equation could be used to find the dimensions of the rug? a. L(4-L) = 120 b. L(L-4) = 120 c. 2(L-4) + 2L = 120 d. 2(4-L) + 2L = 120

OpenStudy (abbycross167):

@Fifciol do you think you could try to help me please?

OpenStudy (xguardians):

C, because a rectangle has 4 sides so you need 4 measurements. 2 L and 2 W W = (L-4) Therefore the equation is: 2(L-4) + 2L = 120

OpenStudy (abbycross167):

Thank you very much!!!